OpenStudy (anonymous):

Find out if the function f(x) is invertible in (0,+inf). In affirmative case, find the inverse function.

OpenStudy (anonymous):

what is the function

OpenStudy (anonymous):



OpenStudy (anonymous):

can you show the work, please

OpenStudy (anonymous):

one way to show that it is invertible is to find the inverse and see if you have a function. another way is to graph and see if it passes "horizontal line test" and a third way is to write \[f(a)=f(b)\] and see if you can conclude that \[a=b\]

OpenStudy (anonymous):

this function says 1) cube 2) add 1 3) exponentiate we can undo that by 1) take the log 2) subtract 1 3) take the cube root so it is invertible and the inverse is \[f^{-1}(x)=\sqrt[3]{\ln(x)-1}\]

OpenStudy (anonymous):

if you want to "show work" you can write \[x=e^{y^3+1}\] and solve for y. you will get the same thing

OpenStudy (anonymous):

Thanks